OpenStudy (anonymous):

Find n such that the error in the approximation of the definite integral is less than .00001 using the trapezoidal rule.

OpenStudy (anonymous):

\[\int\limits_{0}^{1}1\div(1+x)dx\]

OpenStudy (anonymous):

and another: \[\int\limits_{0}^{\pi/2} sinx dx\]